OpenStudy (anonymous):

HighLevel math: find the measure of the the acute angle between the lines 2x+y = 5 and 3x-2y = 8

OpenStudy (amistre64):

create vectors from your slope to play with

OpenStudy (amistre64):

a=<2,-1> and b=<2,3> cos(t) = a.b/|a||b|

OpenStudy (anonymous):

I still don't get the expansion

OpenStudy (amistre64):

\[cos(t)=\frac{2(2)+-1(3)}{\sqrt{(5)(13)}}\] \[t=cos^{-1}\left(\frac{2(2)+-1(3)}{\sqrt{(5)(13)}}\right)\]

OpenStudy (amistre64):

my "a" is a litle off; I got the components swapped about i believe

OpenStudy (amistre64):

2x+y = 5 -2/1 = y/x ... <1,-2> 3x-2y = 8 -3/-2 = y/x ... <2,3>

OpenStudy (amistre64):

yeah, .... so to adjust for errors lol \[cos(t)=\frac{1(2)-2(3)}{\sqrt{(5)(13)}}\] \[t=cos^{-1}\left(\frac{1(2)-2(3)}{\sqrt{(5)(13)}}\right)\] looks a little better

OpenStudy (anonymous):

Thanks a lot

OpenStudy (anonymous):

emm isn the angle supposed to be acute? because you got an obtuse angle